OpenStudy (anonymous):

It's a word problem and has more than one step: a. The sales representative tells you they also have two floor plans available, but they only have 58 homes still for sale. Write an equation that illustrates the situation. Use x and y to denote floor plan one and floor plan two respectively. x + y = 58 b. The sales representative tells you that floor plan one sells for $175,000 and floor plan two sells for $200,000. She also mentions that all the available houses combined are worth $10,200,000. Write an equation that illustrates this situation. Use the same variables you used in

OpenStudy (anonymous):

x175,000 + y 200,000=10,200,000

OpenStudy (anonymous):

I did that but the problem is solving it I tried figuring out how many of each floor plan with no answer working right

OpenStudy (anonymous):

you can solve for it x175,000 + y 200,000=10,200,000 x + y = 58 system of equation , use substitution method

OpenStudy (anonymous):

y=58-x now every where there is y in first equation replace it with 58-x

OpenStudy (anonymous):

but when you distribute, you come up with a negative x variable

OpenStudy (anonymous):

ok i think i got it

OpenStudy (anonymous):

I got x =56 y=2

OpenStudy (anonymous):

yeah.. im checking it now

OpenStudy (anonymous):

omg that works

OpenStudy (anonymous):

thank you very much